HUBBLE SPACE TELESCOPE DAILY REPORT #2543 PERIOD COVERED: 0000Z (UTC) 01/19/00 - 0000Z (UTC) 01/20/00 Daily Status Report as of 020/0000Z 1.0 OBSERVATIONS SCHEDULED AND ACCOMPLISHED: 1.1 Completed Two Sets of STIS/CCD 8408 (STIS CCD Dark Monitor C8) The Space Telescope Imaging Spectrograph (CCD) was used to make a CCD bias calibration observation. This is to monitor the darks for the STIS CCD. The observations were executed as scheduled, and no anomalies were reported. 1.2 Completed TSTIS/CCD 8409 (CCD Bias Monitor C8) The Space Telescope Imaging Spectrograph (CCD) was used to monitor the bias in the 1x1, 1x2, 2x1, and 2x2 bin settings at gain=1, and 1x1 at gain = 4 in order to build up high-S/N superbiases and track the evolution of hot columns. The observations completed with no problems. 1.3 Completed STIS/CCD/MA2 8511 (Full-field Sensitivity and Geometry) The Space Telescope Imaging Telescope (CCD and MA2) was used to measure the geometric distortion for each STIS detector, using the same field as previously used for measuring the MAMA geometric distortion in cycle 7. These exposures will also serve to measure the image location for each of the MAMA imaging modes, with respect to the CCD frame, In order to fix the detector coordinates for the STIS MAMA imaging modes with respect to the CCD coordinates, and hence to the FGS system. The proposal completed nominally. 1.4 Completed STIS/CCD 8202 (Bright Quasar Close Lensing Search) The Space Telescope Imaging Spectrograph (CCD) was used to make observations of HCLQ2316-0 as part of a second generation survey to search for bright lensing quasars. With the restored PSF, new images will be considerably more sensitive to the presence of both close separation and faint lens components, and because STIS reaches two magnitudes fainter the PC, the new snapshot images will have much greater dynamic range. The observations were executed as scheduled, and no anomalies were noted. 1.5 Completed STIS/CCD/MA1/MA2 8424 (MAMA Sensitivity and Focus Monitor C8) The Space Telescope Imaging Spectrograph (CCD, MA1 and MA2) was used to make observations of object J0307+105 in order to measure the sensitivity and to focus monitor C8. This also is to monitor sensitivity of each MAMA grating mode to detect any change due to contamination or other causes, and to monitor the STIS focus in a spectroscopic mode. A wavelength calibration was also taken. The observations were completed as planned, and no anomalies were reported. 1.6 Completed STIS/MA2 8509 (NUV MAMA Darks) The Space Telescope Imaging Spectrograph (MA2) was used to determine the health of the FUV MAMA detector by monitoring the background count rate. The observations completed with no reported problems. 1.7 Completed STIS/CCD 8474 (Coronographic Imaging of Protoplanetary Disks) The Space Telescope Imaging Spectrograph (CCD) was used to obtain broad-band optical coronographic imaging of HD-36112 that will be used to search for circumstellar disks and indirect signatures of the presence of massive planets toward the nearby Herbig Ae/Be and T Tauri stars. The observations completed with no reported problems. 1.8 Completed FGS 8522 (Adjusting FGS2r's AMA for Optimal Interferometric Performance) The FGS2R was used to obtain S-curves at selected locations within the FGS2R FOV. These data will be will be used to adjust the AMA so as to align the HST pupil onto the interferometer's Koesters prisms. The final AMA position will be set to produce an acceptable set of s-curves across the FGS2r FOV. The proposal completed with no reported problems. 1.9 Completed WF/PC-2 8194 (Evolution of High-Redshift Seyfert Galaxies) The WF/PC-2 was used to make observations of the high-redshifted Seyfert galaxy Gal-105016 in order to study its evolution. The observations will provide insights into the Seyfert phenomenon at intermediate redshift and will complement similar HST studies of normal galaxies, clusters and quasars. The observations were executed as scheduled, and no anomalies were noted. 2.0 FLIGHT OPERATIONS SUMMARY: 2.1 Guide Star Acquisitions: Scheduled Acquisitions: 9 Successful: 9 Scheduled Re-acquisitions: 5 Successful: 5 2.2 FHST Updates: Scheduled: 20 Successful: 20 2.3 Operations Notes: Servicing Mission Observatory Verification (SMOV) activities comprise much of the operational activity. Using ROP SR-1, the SSR EDAC error counter was cleared once. SSA transmitters 1 and 2 were turned on and off once each per ROP IC-1. The same activities were accomplished for MA transmitter 1 per ROP IC-2. The STIS MCE-1 reset at 019/142602Z while the low voltage was on and while in an SAA interval. The STIS flight software error counter was cleared at 019/1453Z. MAMA-1 will be recovered at 020200934Z via normal SMS commanding. 3.0 SIGNIFICANT FORTHCOMING EVENTS: Continuation of SMOV and normal science observations and calibrations. /CAW